### v1

# Fuzzy Aristotelian Diagrams

Abstract
I am describing the square of opposition, in particular, and, Aristotelian Diagrams, in general. Then I describe how one can create a mathematical universe to host them. Based on this work, I introduce fuzzy Aristotelian Diagrams and describe a mathematical formulation of them. In addition, I outline the cdharacteristrics of a mathematical universe that can host them.
1 Introduction
The square of opposition is a diagram that graphically represents a collection of logical relationships [9]. This diagram was particularly useful for logicians. More specifically, they used it in their for work for more than 2000 years. The ideas behind the square of opposition have their roots in the work of Aristotle, who lived in the fourth century BC. The square of opposition is based on the rule of contradictory pairs that is discussed in De Interpretatione 6–9. This rule states that for any contradictory pair of statements, only one is true and the other is false. De Interpretatione or On Interpretation (Greek: Per’i ¡ermhne’ias; Latinized Greek: Peri Hermeneias) is the second book of Aristotle’s Organon and deals with the relationship between language and logic [5]. As it happens with many ideas and scientific results of the past, the square of opposition has been severely criticized in recent decades, nevertheless, it is still something that is frequently used by logicians.
The underlying logic of the square of oppostition is the classical bivalent logic. Some call it Aristotelian logic, although in chapter 9 of De Interpretatione, Aristotle asked whether it makes sense to say that a sentence about a future event that can occur or cannot occur is true or false. Thus, the problem of future contingents, as it is called, indirectly speculated about more than two truth values.
Fuzzy sets and their extensions (i.e., Krassimir Atanassov’s intuitionistic fuzzy sets [1]) are mathematical tools that are used to deal with vagueness. But what is vagueness? Generally speaking, as David Lanius [6] correctly points out, vagueness is something that leads to the Sorites paradox, borderline cases, and the violation of the logical principle of bivalent logic [readers not familiar with the Sorites paradox and borderline cases should check out any presentation in layman’s terms (e.g., see [14])]. However, Lanius claims that vagueness is something bad! In my own opinion, this is wrong since vagueness is a fundamental property of our world. Currently, quantum mechanics is our best theory that accurately describes the building blocks of our universe and this theory is vague theory! Why? Because vagueness is the underlying property that affects the properties of the building blocks of our cosmos (see [16, 13, 15]). Of course, vagueness is a semantic property of linguistic expressions [2], but this is something I will not further explore.
Probability theory is a branch of mathematics that studies the outcome of random experiments (e.g., the roll of a die, the results of horse racing, the maximum and the minimum temperature in New York in two weeks from today, etc.). For any such random experiment we need to be aware of all possible outcomes and then to see whether the outcome of such an experiment is really random (i.e., we cannot predict it). Depending on the characteristics of the experiment it is possible to assign probabilities to the possible outcomes. These probabilities are numbers that belong to the unit interval. For example, when we throw a balanced, six-faced die the possible outcomes are 1, 2, 3, 4, 5, 6 and to each of them we assign a probability of $1/6$. Thus, the probability that the next roll of the die will be 6 is $1/6$. Some authors consider that probabilities and membership degrees are the same thing or at least two facets of the same concept. In fact, Zadeh, the founder of fuzzy set theory, believed this. However, this idea is totally wrong from a pragmatic and, more generally, a philosophic point of view. Suffice it to say that probabilities and membership degrees measure completely dif and only iferent things. A probability is the likelihood degree of some event while a membership degree is a truth value (i.e., something that asserts to what degree something has a particular property or what degree somerthing will happen). This is exactly the rerason why I do not find useful the idea of proposing a probabilistic version of the square of opposition [10].
It is my firm belief that a reinterpetation of the square of opposition using fuzzy mathematics is quite reasonable. For examle, one such work uses functional degrees of inclusion to reinterpretate contradiction [7]. Howerver, I am not merely interested in fuzzifying the square of opposition, but to make it “alive” in a vague universe. Such a universe could be a vague category [8, 12]. Leander Vignero [20] has already examined some properties of a category whose object are squares of opposition. In addition, Alexander De Klerck, Leander Vignero, and Lorenz Demey [4] defined dif and only iferent morphisms between squares of opposition and thus defined dif and only iferent categories.
Plan of the chapter
First I will properly introduce the square of opposition. Then, I will explain how one can fuzzify it using intuitionistic fuzzy logic, and then I will show how to build a proper universe for them. The chapter concludes with the customary cinclusions section.
2 Traditional Aristotelian Diagrams
The square of opposition looks like a commutative diagram, however, its corners are logical forms. The table that follows describes the four logical forms.
FORM TITLE
Every $S$ is $P$ Universal Affirmative
No $S$ is $P$ Universal Negative
Some $S$ is $P$ Particular Affirmative
Some $S$ is not $P$ Particular Negative
Figure 1 shows the traditional square of opposition.
Figure 1: The traditional square of opposition.
The important question is: What is the meaning of this diagram? In dif and only iferent words, which facts are conveyed by this diagram? The following theses, as the they are called by the author of [9], are expressed in diagram 1.
  • “Every $S$ is $P$” and “Some $S$ is not $P$” are contradictories.
  • “No $S$ is $P$” and “Some $S$ is $P$” are contradictories.
  • “Every $S$ is $P$ and “No $S$ is $P$ are contraries.
  • “Some $S$ is $P$” and “Some $S$ is not $P$ are subcontraries.
  • “Some $S$ is $P$” is a subaltern of “Every $S$ is $P$.”
  • “Some $S$ is not $P$’ is a subaltern of “No $S$ is $P$.”
A modern version of the square of opposition is shown in figure 2. Since this diagram conveys little information, it is not used frequently.
Figure 2: A modern version of the square of opposition.
Mathematically speaking, a square of opposition is an Aristotelian diagram. These diagrams are described by Boolean algebras. But, what is a Boolean algebra? The following definition explains what is a Boolean algebra. [^1] Weisstein, Eric W. “Boolean Algebra.” From MathWorld–A Wolfram Web Resource. https://mathworld.wolfram.com/BooleanAlgebra.html
Definition 2.1
A Boolean algebra is a set $B$ of elements $a,b,\ldots$ with the following properties:
  1. 1.
    $B$ is equipped with two binary operations, $\wedge$ and $\vee$, which satisfy the idempotent laws
    $a\wedge a=a\vee a=a,$
    the commutative laws
    $a\wedge b$ $=b\wedge a$
    $a\vee b$ $=b\vee a,$
    and the associative laws
    $a\wedge(b\wedge c)$ $=(a\wedge b)\wedge c$
    $a\vee(b\vee c)$ $=(a\vee b)\vee c.$
  2. 2.
    In addition, these operations satisfy the absorption law
    $a\wedge(a\vee b)=a\vee(a\wedge b)=a.$
  3. 3.
    Also, the operations are mutually distributive
    $a\wedge(b\vee c)$ $=(a\wedge b)\vee(a\wedge c)$
    $a\vee(b\wedge c)$ $=(a\vee b)\wedge(a\vee c).$
  4. 4.
    $B$ contains universal bounds 00 and $1$ which satisfy
    $0\wedge a$ $=0$
    $0\vee a$ $=a$
    $1\wedge a$ $=a$
    $1\vee a$ $=1.$
  5. 5.
    $B$ has a unary operation $\neg a$ of complementation, which obeys the laws
    $a\wedge\neg a=0\quad\text{and}\quad a\vee\neg a=1.$
From here we can easily define Aristotelian diagrams [4]:
Definition 2.2
An Aristotelian diagram $D$ is a pair $(F,B)$, where $B$ is a Boolean algebra $(B,\wedge_{B},\vee_{B},\neg_{B},1_{B},0_{B})$ and $F\subseteq B$. When the Boolean algebra $B$ is clear from context, we usually omit the subscripts from $\wedge$, $\vee$, etc.
The next thing we need to know is how the theses described above are expressed in the language of Boolean algebras.
Definition 2.3
Assume that $(F,B)$ is an Aristolenian diagram. Then, we say that $x,y\in B$ are:
  • $B$-bi-implication ($\text{BI}_{B}$) if and only if $x=y$;
  • $B$-left-implication ($\text{LI}_{B}$) if and only if $x<y$;
  • $B$-right-implication ($\text{RI}_{B}$) if and only if $y<x$;
  • $B$-contradictory ($\text{CD}_{B}$) if and only if $x\mathbin{\wedge_{B}}y=0_{B}$ and $x\mathbin{\vee_{B}}y=1_{B}$, that is, $x=\mathop{\neg_{B}}y$;
  • $B$-contrary ($\text{C}_{B}$) if and only if $x\mathbin{\wedge_{B}}y=0_{B}$ and $x\mathbin{\vee_{B}}y\neq 1_{B}$, that is, $x\mathrel{<_{B}}\mathop{\neg_{B}}y$;
  • $B$-subcontrary ($\text{SC}_{B}$) if and only if $x\mathbin{\wedge_{B}}y\neq 0_{B}$ and $x\mathbin{\vee_{B}}y=1_{B}$, that is, that is, $x\mathrel{>_{B}}\mathop{\neg_{B}}y$;
  • $B$-unconnectedness ($\text{Un}_{B}$) if and only if none of the above holds.
The above relations are called logical relations and are denoted by $\Re$.
Definition 2.4
An Aristotelian isomorphism $f:(F_{1},B_{1})\rightarrow(F_{2},B_{2})$ is a bijection $f:F_{1}\rightarrow F_{2}$ such that for all logical relations $R\in\Re$ and all $x,y\in F_{1}$ we have that $x\mathrel{R_{B_{1}}}y$ if and only if $f(x)\mathrel{R_{B_{2}}}f(y)$.
This is strong way to go from one diagram to another and is not bery useful. A weaker one make use of the following order relation:
Definition 2.5
There is an informativity order $\leq_{i}$ on $\Re$ which is given by: $\text{Un}\mathrel{\leq_{i}}\text{LI}$, $\text{Un}\mathrel{\leq_{i}}\text{RI}$, $\text{Un}\mathrel{\leq_{i}}\text{C}$, $\text{Un}\mathrel{\leq_{i}}\text{SC}$, $\text{LI}\mathrel{\leq_{i}}\text{BI}$, $\text{RI}\mathrel{\leq_{i}}\text{BI}$, $\text{C}\mathrel{\leq_{i}}\text{CD}$, and $\text{SC}\mathrel{\leq_{i}}\text{CD}$.
Infomorphisms are defined as follows:
Definition 2.6
An infomorphism $f:(F_{1},B_{1})\rightarrow(F_{2},B_{2})$ is a function that satisfies the following condition: for all $x,y\in F_{1}$ it holds that if $x\mathrel{R_{B_{1}}}y$, then $f(x)\mathrel{S_{B_{2}}}f(y)$ with $R\mathrel{\leq_{i}}S$.
3 Contradiction and Vagueness
If we want to interpret diagram 1 using a fuzzy logic (i.e., a logic of vagueness), then we must understand how contradiction is represented in such a logic. Trillas, Alsina, and Jacas [18] started from the classical law of contradiction, which says that for all propositions $p$, it is impossible for both $p$ and not $p$ to be true, that is, $p\Rightarrow\neg p$. Such propositions are called self-contradictory. In addition, if and only if $p\Rightarrow\neg q$ is true, then $p$ and $q$ are contradictory. However, I have to remark that the standard law of contradiction states that for all $p$, $p\wedge\neg p$ is always false. Trillas et al., assume that in general $p\Rightarrow q\equiv\neg p\vee(p\wedge q)$, that is, the implication operator is the so-called material impication. If we assume that $p$ and $q$ are fuzzy sets (i.e., the equivalent of predicates in classical logic), then we can define the degree of contradiction as follows:
Definition 3.1
Suppose that $A,B\in[0,1]^{E}$, and $J$ is an implication operator, and $N$ a negation operator, then $A$ is contradictory to $B$ with degree equal to $r\in[0,1]$, when $J(A(x),N(B(x))=r$.
From here we define distances and other things (e.g., see [3, 17]).
Assume that we assign truth values to vague propositions that belong to the unit interval, nevertheless, we do not follow the principles of fuzzy mathematics. Indeed, Nicholas Smith [11] went to great lengths to propose that contradictory propositions should always have a truth value that is equal to $0.5$. In my own opionion it is more natural to accept Smith’s proposal and here is why. Assume that an eagle flies in the sky and consider the statement “the eagle is in a cloud.” According to classical logic this statement can be either true or false. However, when the eagle is partly inside a cloud and, consequently, partly outside the cloud, then, strictly speaking, the statement is false but, it is closer to reality to assume that the statement and its negation are both true. After all, this why paraconsistent logics (i.e., logics that assume that some contradictions are true) are considered bivalent models of vagueness. In conclusion, it makes sense to assume that the degree to which a vague proposition is a contradiction is always $0.5$. Obviously, this means that the degree to which it is not a contradiction should be $0.5$. I think it makes sense that this degree is less than or equal to $0.5$. In addition, this value should not be derivable from some formula. In dif and only iferent words, I propose that “intuitionistic” fuzzy logic (IFL) is the ideal tool to model vague contradictions. Figure 3 shows how the traditional square of opposition can be modified when the underlying logic is IFL.
Figure 3: An “intuitionistic” fuzzy version of the traditional square of opposition.
4 “Intuitionistic” Fuzzy Aristotelian Diagrams
“Intuitionistic” fuzzy Aristotelian diagrams are subsets of “intuitionistic” fuzzy Boolean algebras. In order to properly define these structures, we need some auxiliary definitions stated in [19].
Definition 4.1
An “intuitionistic” fuzzy relation is an “intuitionistic” fuzzy subset of $X\times Y$; that is, the set $R$ given by
$R=\bigl{\{}\langle(x,y),\mu_{R}(x,y),\nu_{R}(x,y)\rangle\mathrel{\bigm{|}}x\in X,y\in Y\bigr{\}},$
where $\mu_{R},\nu_{R}:X\times Y\rightarrow[0,1]$, satisfy the condition $0\leq\mu_{R}(x,y)+\nu_{R}(x,y)\leq 1$, for every $(x,y)\in X\times Y$.
Definition 4.2
An“intuitionistic” fuzzy relation in $X\times X$ is
reflexive
if for every $x\in X$, $\mu_{R}(x,x)=1$ and $\nu_{R}(x,x)=0$.
perfectly antisymmetric
if for every $(x,y)\in X\times X$ with $x\neq y$ and $\mu_{R}(x,y)>0$ or $\mu_{R}(x,y)=0$ and at the same time $\nu_{R}(x,y)<1$, then $\mu_{R}(x,y)=0$ and $\nu_{R}(x,y)=1$.
transitive
if $R\circ R\subseteq R$, where $\circ$ is $\max$-$\min$ and $\min$-$\max$ composition, that is
$\mu_{R}(x,z)\geq\max_{y}\Bigl{[}\min\bigl{\{}\mu_{R}(x,y),\mu_{R}(x,y)\bigr{\}}\Bigr{]}$
and
$\nu_{R}(x,z)\leq\min_{y}\Bigl{[}\max\bigl{\{}\nu_{R}(x,y),\nu_{R}(y,z)\bigr{\}}\Bigr{]}.$
Definition 4.3
An “intuitionistic” fuzzy relation $R$ on $X\times X$ is said to be an “intuitionistic” fuzzy partially ordered relation if $R$ is reflexive, perfectly antisymmetric, and transitive.
Definition 4.4
A crisp set $X$ on which an “intuitionistic” fuzzy partial ordering $R$ is defined is said to be an “intuitionistic” fuzzy lattice if and only if for any two element set $\{x,y\}\subset X$, the least upper bound (lub) and greatest lower bound (glb) exist in $X$. We denote the lub of $\{x,y\}$ by $x\vee y$ and the glb of $\{x,y\}$ by $x\wedge y$.
Definition 4.5
An “intuitionistic” fuzzy lattice $(X,R)$ is distributive if and only if for all $a,b,c\in X$,
$a\wedge(b\vee c)=(a\wedge b)\vee(a\wedge c)\quad\text{and}\quad a\vee(b\wedge c)=(a\vee b)\wedge(a\vee c).$
Theorem 4.1
If $(X,R)$ is a complemented (i.e., it has a unary operation $\neg a$ of complementation), distributive “intuitionistic” fuzzy lattice, then the two De Morgan’s Laws
$\neg(a\vee b)=\neg a\wedge\neg b\quad\text{and}\quad\neg(a\wedge b)=\neg a\vee\neg b,$
hold for all $a,b\in X$.
Obviously, we can use the standard complementaion of “intuitionistic” fuzzy sets or a variation of it suitable for our case.
Definition 4.6
A complemented distributive ‘intuitionistic” fuzzy lattice is an ‘intuitionistic” fuzzy Boolean algebra.
And now comes the definition of fuzzy Aristotelian diagrams:
Definition 4.7
Assume that $B=(X,R)$ is an “intuitionistic” fuzzy Boolean algebra. Then, a fuzzy Aristotelian diagram $D$ is a pair $(F,B)$, where $F\subseteq X$.
5 A Fuzzy Category of Fuzzy Aristotelian Diagrams
Categories of fuzzy subsets or other fuzzy structures have been proposed by this author (see [12] and references therein). However, here I am going tpo define a new mathematical universe where these “intuitionistic” fuzzy squares of opposition will live. To do so, I will extend the structures descibed in [12]. This structures are categories but each morphism has a plausibility degree (i.e., a number) that determines the degree to which the domain can “go” to the codomain. In more strict mathematical parlance, this numbner is the degree to which this morphism is plausible. The point of departure is that in what follows I will assume that the underlying logic will be an “intuitionistic” fuzzy logic. The result will be a pure “intuitionistic” fuzzy category. First let me remind the reader what is a category:
Definition 5.1
A category $\mathcal{C}$ is made up of
  1. 1.
    a collection of things that are called $\mathcal{C}$-objects (objects for simplicity);
  2. 2.
    a collection of “bridges” between $\mathcal{C}$-objects that are called $\mathcal{C}$-arrows (arrows fro simplicity) or $\mathcal{C}$-morphisms (or just morphisms for simplicity);
  3. 3.
    each arrow $f$ has as domain the object $\mathop{\text{dom}}f$ and as codomain the object $\mathop{\text{cod}}f$. If $A=\mathop{\text{dom}}f$ and $B=\mathop{\text{cod}}f$, then we write $f:A\rightarrow B$;
  4. 4.
    an operation that assigns to each pair $(g,f)$ of $\mathcal{C}$-arrows, such that $\mathop{\text{dom}}g=\mathop{\text{cod}}f$, a $\mathcal{C}$-arrow $g\circ f$, such that $\mathop{\text{dom}}(g\circ f)=\mathop{\text{dom}}f$ and $\mathop{\text{cod}}(g\circ f)=\mathop{\text{cod}}g$, that is, $g\circ f:\mathop{\text{dom}}f\rightarrow\mathop{\text{cod}}g$. In addition, given the arrows $f:A\rightarrow B$, $g:B\rightarrow C$, and $h:C\rightarrow D$, then $h\circ(g\circ f)=(h\circ g)\circ f$;
  5. 5.
    for each $\mathcal{C}$-object $A$ there is a $\mathcal{C}$-arrow $\text{id}_{A}:A\rightarrow A$ called the identity arrow, such that for any $f:A\rightarrow B$ and $g:B\rightarrow C$, $\text{id}_{B}\circ f=f$ and $g\circ\text{id}_{B}=g$.
Let me now define fuzzy categories:
Definition 5.2
A fuzzy category $\mathscr{C}$ is an ordinary category $\mathcal{C}$ but in addition:
  1. 1.
    There is an operation $\mathop{\mathrm{p}}$ that assigns to each arrow a plausibility degree $\rho=\mathop{\mathrm{p}}f\in[0,1]$. Thus, an arrow that starts from $A$ and ends at $B$ with plausibility degree $\rho$ is written as:
[…]
  2. 2.
    For the composite $g\circ f$ it holds that $\mathop{\mathrm{p}}(g\circ f)=(\mathop{\mathrm{p}}f)\wedge(\mathop{\mathrm{p}}g)$. The associative law holds since $\wedge$ is an associative operation.
  3. 3.
[…]
    $\mathbf{1}_{B}\circ f=f\quad\mbox{and}\quad g\circ\mathbf{1}_{B}=g$
    for any $\mathscr{C}$-arrows $f:A\overset{\rho_{f}}{\longrightarrow}B$ and $g:B\overset{\rho_{g}}{\longrightarrow}A$.
To define “intuitionsistic” fuzzy categories I need to make only a small adjustment to the previous definition:
Definition 5.3
A “intuitionsistic” fuzzy category $\mathscr{C}$ is a fuzzy category $\mathscr{C^{\prime}}$ but in addition there is an operation $\mathop{\overline{\mathrm{p}}}$ that assigns to each arrow a non-plausibility degree $\sigma=\mathop{\overline{\mathrm{p}}}f\in[0,1]$. Thus, an arrow that starts from $A$ and ends at $B$ with plausibility degree $\rho$ and non-plausbility degree $\sigma$ is written as:
[…]
Now it is possible to define a universe of fuzzy Aristotelian diagrams. However, it is necessary to define morphisms between them. Surprisingly, we can use something similar to the morphisms described in section 2. However, a detailed description of these morphisms is not available yet.
6 Conclusions
I have described (crisp) Aristotelian diagrams and I introduced a fuzzy version of these diagrams. In addition, I described the general characteristics of a mathematical universe that can host them. The mechanism to go from one diagram to another and, consequently, a method to compose them is a subject of actrive research now.