### v2

# Fuzzy Aristotelian Diagrams

Abstract
After a concise introduction to the square of opposition, in particular, and, Aristotelian Diagrams, in general, I describe how one can create a mathematical universe to host these objects. Since these objects assume that the underlying logic is the bivalent logic, I have used these objects as a starting point to introduce fuzzy Aristotelian Diagrams and describe a mathematical formulation of them. In addition, I outline the cdharacteristrics of a mathematical universe that hosts them.
1 Introduction
The square of opposition is a diagram that graphically represents a collection of logical relationships [11]. This diagram was particularly useful for logicians. More specifically, they used it in their work for more than 2000 years. The ideas behind the square of opposition have their roots in the work of the Greek philosopher Aristotle, who lived in the fourth century BC. De Interpretatione 6–9 discusses the rule of contradictory pairs, which forms the basis of the square of opposition. This rule states that for any contradictory pair of statements, only one is true and the other is false. De Interpretatione or On Interpretation (Greek: Per’i ¡ermhne’ias; Latinized Greek: Peri Hermeneias) is the second book of Aristotle’s Organon and deals with the relationship between language and logic [7]. As it happens with many ideas and scientific results of the past, the square of opposition has been severely criticized in recent decades; nevertheless, it is still something that is frequently used by logicians.
The underlying logic of the square of opposition is the classical bivalent logic. Some call it Aristotelian logic, although in chapter 9 of De Interpretatione, Aristotle asked whether it makes sense to say that a sentence about a future event that can occur or cannot occur is true or false. Thus, the so-called problem of future contingents indirectly speculates about more than two truth values.
We use fuzzy sets and their extensions, such as Krassimir Atanassov’s intuitionistic fuzzy sets [1], as mathematical tools to address vagueness. But what is vagueness? Generally speaking, David Lanius [8] correctly points out that vagueness leads to the Sorites paradox, borderline cases, and the violation of the logical principle of bivalent logic; readers who are not familiar with the Sorites paradox and borderline cases should consult any presentation in layman’s terms (e.g., see [16]). However, Lanius claims that vagueness is inherently negative. In my own opinion, this claim is wrong since vagueness is a fundamental property of our world. Currently, quantum mechanics is our best theory for accurately describing the building blocks of our universe, and it is considered a vague theory! Why? Because the properties of our cosmos’ building blocks are influenced by the underlying property of vagueness (refer to [19, 15, 17]). Of course, vagueness is a semantic property of linguistic expressions [3], but this is something I will not further explore.
Probability theory is a branch of mathematics that studies the outcome of random experiments (e.g., the roll of a die, the results of horse racing, the maximum and the minimum temperature in New York in two weeks from today, etc.). For any random experiment, we must be aware of all possible outcomes and determine whether the result of the experiment is truly random (i.e., unpredictable). Depending on the characteristics of the experiment, probabilities can be assigned to the possible outcomes. These probabilities are numbers that belong to the unit interval. For example, when we throw a balanced, six-faced die, the possible outcomes are 1, 2, 3, 4, 5, and 6, and to each of them we assign a probability of 1/6. Thus, the probability that the next roll of the die will be 6 is 1/6. Some authors consider that probabilities and membership degrees are the same thing or at least two facets of the same concept. In fact, Zadeh, the founder of fuzzy set theory, believed this. However, this idea is totally wrong from a pragmatic and, more generally, a philosophic perspective. It’s important to note that probabilities and membership degrees measure entirely distinct entities. A probability is the likelihood degree of some event, while a membership degree is a truth value (i.e., something that asserts to what degree something has a particular property or to what degree something will happen). This is precisely why I do not find the idea of proposing a probabilistic version of the square of opposition [12] to be useful.
It is my firm belief that a reinterpretation of the square of opposition using fuzzy mathematics is quite reasonable. For example, one such work uses functional degrees of inclusion to reinterpret contradiction [9]. However, I am not merely interested in fuzzifying the square of opposition but in making it “alive” in a vague universe. Such a universe could be a vague category [10]. Leander Vignero [23] has already examined some properties of a category whose objects are squares of opposition. In addition, Alexander De Klerck, Leander Vignero, and Lorenz Demey [5] defined different morphisms between squares of opposition and thus defined different categories.
Plan of the paper
First I will properly introduce the square of opposition. Next, I will explain how to fuzzify the square of opposition using intuitionistic fuzzy logic, and afterward, I will demonstrate how to construct a suitable universe for these concepts. The paper concludes with the customary conclusions section.
2 Traditional Aristotelian Diagrams
The square of opposition looks surprisingly like a commutative diagram (see Figure 1); however, its corners are logical forms and not some ordinary objects. The table that follows describes these four logical forms.
FORM TITLE
Every $S$ is $P$ Universal Affirmative
No $S$ is $P$ Universal Negative
Some $S$ is $P$ Particular Affirmative
Some $S$ is not $P$ Particular Negative
Figure 1: The traditional square of opposition.
If the diagram in Figure 1 is not a commutative diagram, then what is the meaning of this diagram? Put differently, what facts does this diagram convey? The following theses, as they are called by the author of [11], are expressed in the diagram of Figure 1.
  • “Every $S$ is $P$” and “Some $S$ is not $P$” are contradictories.
  • “No $S$ is $P$” and “Some $S$ is $P$” are contradictories.
  • “Every $S$ is $P$ and “No $S$ is $P$ are contraries.
  • “Some $S$ is $P$” and “Some $S$ is not $P$ are subcontraries.
  • “Some $S$ is $P$” is a subaltern of “Every $S$ is $P$.”
  • “Some $S$ is not $P$’ is a subaltern of “No $S$ is $P$.”
Figure 2 shows a modern version of the square of opposition. People rarely use this diagram because it conveys little information.
Figure 2: A modern version of the square of opposition.
Mathematically speaking, a square of opposition is an Aristotelian diagram and Boolean algebras describe these diagrams. But what is a Boolean algebra? The following definition explains what a Boolean algebra is. [^1] Weisstein, Eric W. “Boolean Algebra.” From MathWorld–A Wolfram Web Resource. https://mathworld.wolfram.com/BooleanAlgebra.html
Definition 2.1
A Boolean algebra is a set $B$ of elements $a,b,\ldots$ with the following properties:
  1. 1.
    $B$ is equipped with two binary operations, $\wedge$ and $\vee$, which satisfy the idempotent laws
    $a\wedge a=a\vee a=a,$
    the commutative laws
    $a\wedge b$ $=b\wedge a$
    $a\vee b$ $=b\vee a,$
    and the associative laws
    $a\wedge(b\wedge c)$ $=(a\wedge b)\wedge c$
    $a\vee(b\vee c)$ $=(a\vee b)\vee c.$
  2. 2.
    In addition, these operations satisfy the absorption law
    $a\wedge(a\vee b)=a\vee(a\wedge b)=a.$
  3. 3.
    Also, the operations are mutually distributive
    $a\wedge(b\vee c)$ $=(a\wedge b)\vee(a\wedge c)$
    $a\vee(b\wedge c)$ $=(a\vee b)\wedge(a\vee c).$
  4. 4.
    $B$ contains universal bounds 00 and $1$ which satisfy
    $0\wedge a$ $=0$
    $0\vee a$ $=a$
    $1\wedge a$ $=a$
    $1\vee a$ $=1.$
  5. 5.
    $B$ has a unary operation $\neg a$ of complementation, which obeys the following laws:
    $a\wedge\neg a=0\quad\text{and}\quad a\vee\neg a=1.$
Using this defintion, we define Aristotelian diagrams [5]:
Definition 2.2
An Aristotelian diagram $D$ is a pair $(F,B)$, where $B$ is the underlying set of the Boolean algebra $(B,\wedge_{B},\vee_{B},\neg_{B},1_{B},0_{B})$ and $F\subseteq B$. When the Boolean algebra is clear from context, we usually omit the subscripts from $\wedge$, $\vee$, etc.
The next thing we need to know is how the theses described above are expressed in the language of Boolean algebras.
Definition 2.3
Assume that $(F,B)$ is an Aristolenian diagram. Then, we say that $x,y\in B$ are in:
  1. 1.
    $B$-bi-implication ($\text{BI}_{B}$) if and only if $x=y$;
  2. 2.
    $B$-left-implication ($\text{LI}_{B}$) if and only if $x<y$;
  3. 3.
    $B$-right-implication ($\text{RI}_{B}$) if and only if $y<x$;
  4. 4.
    $B$-contradictory ($\text{CD}_{B}$) if and only if $x\mathbin{\wedge_{B}}y=0_{B}$ and $x\mathbin{\vee_{B}}y=1_{B}$, that is, $x=\mathop{\neg_{B}}y$;
  5. 5.
    $B$-contrary ($\text{C}_{B}$) if and only if $x\mathbin{\wedge_{B}}y=0_{B}$ and $x\mathbin{\vee_{B}}y\neq 1_{B}$, that is, $x\mathrel{<_{B}}\mathop{\neg_{B}}y$;
  6. 6.
    $B$-subcontrary ($\text{SC}_{B}$) if and only if $x\mathbin{\wedge_{B}}y\neq 0_{B}$ and $x\mathbin{\vee_{B}}y=1_{B}$, that is, that is, $x\mathrel{>_{B}}\mathop{\neg_{B}}y$;
  7. 7.
    $B$-unconnectedness ($\text{Un}_{B}$) if and only if none of the above holds.
The above relations are called logical relations and are denoted by $\Re$. Also, the first three relations are called implication relations and relations 4–6 are called opposition relations. The following definition provides a formal mechanism for relating Aristolenian diagrams.
Definition 2.4
An Aristotelian isomorphism $f:(F_{1},B_{1})\rightarrow(F_{2},B_{2})$ is a bijection $f:F_{1}\rightarrow F_{2}$ such that for all logical relations $R\in\Re$ and all $x,y\in F_{1}$ we have that $x\mathrel{R_{B_{1}}}y$ if and only if $f(x)\mathrel{R_{B_{2}}}f(y)$.
This is strong way to go from one diagram to another and is not bery useful. A weaker one make use of the following order relation:
Definition 2.5
There is an informativity order $\leq_{i}$ on $\Re$ which is given by: $\text{Un}\mathrel{\leq_{i}}\text{LI}$, $\text{Un}\mathrel{\leq_{i}}\text{RI}$, $\text{Un}\mathrel{\leq_{i}}\text{C}$, $\text{Un}\mathrel{\leq_{i}}\text{SC}$, $\text{LI}\mathrel{\leq_{i}}\text{BI}$, $\text{RI}\mathrel{\leq_{i}}\text{BI}$, $\text{C}\mathrel{\leq_{i}}\text{CD}$, and $\text{SC}\mathrel{\leq_{i}}\text{CD}$.
Infomorphisms are defined as follows:
Definition 2.6
An infomorphism $f:(F_{1},B_{1})\rightarrow(F_{2},B_{2})$ is a function that satisfies the following condition: for all $x,y\in F_{1}$ it holds that if $x\mathrel{R_{B_{1}}}y$, then $f(x)\mathrel{S_{B_{2}}}f(y)$ with $R\mathrel{\leq_{i}}S$.
3 Contradiction and Vagueness
If we want to interpret diagram 1 using fuzzy logic (i.e., a logic of vagueness), then we must understand how contradiction is represented in such a logic. Trillas, Alsina, and Jacas [21] started from the classical law of contradiction, which says that for all propositions $p$, it is impossible for both $p$ and not $p$ to be true, that is, $p\Rightarrow\neg p$. Such propositions are called self-contradictory. In addition, if and only if $p\Rightarrow\neg q$ is true, then $p$ and $q$ are contradictory. However, I have to remark that the standard law of contradiction states that for all $p$, $p\wedge\neg p$ is always false. Trillas et al., assume that in general $p\Rightarrow q\equiv\neg p\vee(p\wedge q)$, that is, the implication operator is the so-called material implication. If we assume that $p$ and $q$ are fuzzy sets (i.e., the equivalent of predicates in classical logic), then we can define the degree of contradiction as follows:
Definition 3.1
Suppose that $A,B\in[0,1]^{E}$, and $J$ is an implication operator, and $N$ a negation operator, then $A$ is contradictory to $B$ with degree equal to $r\in[0,1]$, when $J(A(x),N(B(x))=r$.
From here we can define distances and other things (e.g., see [4, 20]).
Assume that we assign truth values to vague propositions that belong to the unit interval; nevertheless, we do not follow the principles of fuzzy mathematics. Indeed, Nicholas Smith [14] proposed that contradictory propositions should always have a truth value equal to $0.5$. In my opinion, it is more natural to accept Smith’s proposal, and here is why. Assume that an eagle flies in the sky and consider the statement, “The eagle is in a cloud.” According to classical logic, this statement can be either true or false. However, when the eagle is partly inside a cloud and, consequently, partly outside the cloud, then, strictly speaking, the statement is false, but it is closer to reality to assume that the statement and its negation are both true. After all, this is why paraconsistent logics (i.e., logics that assume that some contradictions are true) are considered bivalent models of vagueness. In conclusion, it makes sense to assume that the degree to which a vague proposition is a contradiction is always $0.5$. This means that the degree to which it is not a contradiction should be $0.5$. I think it makes sense that this degree is less than or equal to $0.5$. In addition, this value should not be derivable from some formula. In different words, I propose that “intuitionistic” fuzzy logic (IFL) is the ideal tool to model vague contradictions. Figure 3 shows how the traditional square of opposition can be modified when the underlying logic is IFL.
Figure 3: An “intuitionistic” fuzzy version of the traditional square of opposition.
4 “Intuitionistic” Fuzzy Aristotelian Diagrams
“Intuitionistic” fuzzy Aristotelian diagrams are subsets of “intuitionistic” fuzzy Boolean algebras. To properly define these structures, we need some auxiliary definitions stated in [22].
Definition 4.1
An “intuitionistic” fuzzy relation is an “intuitionistic” fuzzy subset of $X\times Y$; that is, the set $R$ given by
$R=\bigl{\{}\langle(x,y),\mu_{R}(x,y),\nu_{R}(x,y)\rangle\mathrel{\bigm{|}}x\in X,y\in Y\bigr{\}},$
where $\mu_{R},\nu_{R}:X\times Y\rightarrow[0,1]$, satisfy the condition $0\leq\mu_{R}(x,y)+\nu_{R}(x,y)\leq 1$, for every $(x,y)\in X\times Y$.
Definition 4.2
An“intuitionistic” fuzzy relation in $X\times X$ is
reflexive
if for every $x\in X$, $\mu_{R}(x,x)=1$ and $\nu_{R}(x,x)=0$.
perfectly antisymmetric
if for every $(x,y)\in X\times X$ with $x\neq y$ and $\mu_{R}(x,y)>0$ or $\mu_{R}(x,y)=0$ and at the same time $\nu_{R}(x,y)<1$, then $\mu_{R}(x,y)=0$ and $\nu_{R}(x,y)=1$.
transitive
if $R\circ R\subseteq R$, where $\circ$ is $\max$-$\min$ and $\min$-$\max$ composition, that is
$\mu_{R}(x,z)\geq\max_{y}\Bigl{[}\min\bigl{\{}\mu_{R}(x,y),\mu_{R}(x,z)\bigr{\}}\Bigr{]}$
and
$\nu_{R}(x,z)\leq\min_{y}\Bigl{[}\max\bigl{\{}\nu_{R}(x,y),\nu_{R}(y,z)\bigr{\}}\Bigr{]}.$
Definition 4.3
An “intuitionistic” fuzzy relation $R$ on $X\times X$ is said to be an “intuitionistic” fuzzy partially ordered relation if $R$ is reflexive, perfectly antisymmetric, and transitive.
Definition 4.4
A crisp set $X$ on which an “intuitionistic” fuzzy partial ordering $R$ is defined is said to be an “intuitionistic” fuzzy lattice if and only if for any two element set $\{x,y\}\subset X$, the least upper bound (lub) and greatest lower bound (glb) exist in $X$. We denote the lub of $\{x,y\}$ by $x\vee y$ and the glb of $\{x,y\}$ by $x\wedge y$.
Definition 4.5
An “intuitionistic” fuzzy lattice $(X,R)$ is distributive if and only if for all $a,b,c\in X$,
$a\wedge(b\vee c)=(a\wedge b)\vee(a\wedge c)\quad\text{and}\quad a\vee(b\wedge c)=(a\vee b)\wedge(a\vee c).$
Theorem 4.1
If $(X,R)$ is a complemented (i.e., it has a unary operation $\neg a$ of complementation), distributive “intuitionistic” fuzzy lattice, then the two De Morgan’s Laws
$\neg(a\vee b)=\neg a\wedge\neg b\quad\text{and}\quad\neg(a\wedge b)=\neg a\vee\neg b,$
hold for all $a,b\in X$.
Obviously, we can use the standard complementaion of “intuitionistic” fuzzy sets or a variation of it suitable for our case.
Definition 4.6
A complemented distributive ‘intuitionistic” fuzzy lattice is an ‘intuitionistic” fuzzy Boolean algebra.
And now comes the definition of fuzzy Aristotelian diagrams:
Definition 4.7
Assume that $B=(X,R)$ is an “intuitionistic” fuzzy Boolean algebra. Then, a fuzzy Aristotelian diagram $D$ is a pair $(F,B)$, where $F\subseteq X$.
5 A Category of Fuzzy Aristotelian Diagrams
Let’s now focus on how we can define a universe of fuzzy Aristotelian diagrams. First, it is necesssary to define morphisms between “intuitionistic” fuzzy Boolean algebra. In the classical case, a Boolean homomorphism is a function between two Boolean algebras that preserves their algebraic structure. Since, an “intuitionistic” fuzzy Boolean homomorphism is an extension of the classical concept, it is expected that the new definition will be an extension of the classical one. Indeed, here is how we can define these new morphisms:
Definition 5.1
A function $f$ between two “intuitionistic” fuzzy Boolean algebras $B_{1}$ and $B_{2}$ is a homomorphism if it preserves the operations:
  1. 1.
    $f(x\mathrel{\vee_{B_{1}}}y)=f(x)\mathrel{\vee_{B_{2}}}f(y)$,
  2. 2.
    $f(\mathop{\neg_{B_{1}}}x)=\mathop{\neg_{B_{2}}}f(x)$,
  3. 3.
    $f(0_{B_{1}})=0_{B_{2}}$, and
  4. 4.
    $f(1_{B_{1}})=1_{B_{2}}$.
As in the case of (crisp) Aristotelian diagrams, this definition is too strong and not particularly useful. Thus, I propose the following definition:
Definition 5.2
A fuzzy infomorphism $f:(F_{1},B_{1})\rightarrow(F_{2},B_{2})$ is a function from $F_{1}$ to $F_{2}$ that satisfies the following condition: for all $x,y\in F_{1}$ it holds that if $x\mathrel{R_{B_{1}}}y$, then $f(x)\mathrel{S_{B_{2}}}f(y)$ with $R\mathrel{\leq_{i}}S$.
The following statement can be easily proved.
Proposition 5.1
The collection IFBA of all “intuitionistic” fuzzy Boolean algebras and their fuzzy infomorphism is a category.
Proof.
The objects of the category are “intuitionistic” fuzzy Boolean algebras and the morphisms are the fuzzy infomorphism. The min-max composition is associative (see [18, p. 81]) and the max-min composition is associative for three fuzzy relations (see [13]). Therefore, composition of fuzzy infomorphism is associative. The identity morphism is one that leaves an “intuitionistic” fuzzy Boolean algebra as is. Thus, IFBA is a category. ∎
The various relations presented in the definition 2.3 can be easily extended in the our case. For example, bi-implication is meaningful to a degree. In particular, given an “intuitionistic” fuzzy Boolean algebra $(F,B)$, then for $x,y\in F$, we can say that $x=y$ if and only if the membership degree of $x\mathrel{R_{B}}y$ is equal to the nonmembership degree of $x\mathrel{R_{B}}y$. However, since this is too strict, we can say that the difference of the membership and the nonmembership degrees are within a specific range. For example, we can say that the differences can be less that $0.01$.
6 Conclusions
I have described (crisp) Aristotelian diagrams and I introduced a fuzzy version of these diagrams. In addition, I described the general characteristics of a mathematical universe that can host them. The next thing to do is to examine this mathematical universe and reveal its properties.